OpenStudy (anonymous):

Help me learn the steps in solving this problem pleasee Solve 4x2 + 64 = 0

OpenStudy (anonymous):

are you familiar with Quadratic equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well lets try it for ax^2+bx+c=0 \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] here we have 4x^2+64=0 so comparaing with standard quadratic equation a=4 b=0 c=64 put in the above euation and let me know that yoy get for x.

OpenStudy (anonymous):

im not sure but i got